Line l touches the circle ω at point K. Points A and B are chosen on ω in such a way that they are situated on opposite sides with respect to the diameter of ω that passes through point K. Find the area of triangle AKB if the distances from points A and B to line l are equal to 7 and 11 respectively, and AK=13. Round the answer to the closest integer.
1
Expert's answer
2020-11-26T19:11:30-0500

Let "F\in l, AF\perp l; T\in l, BT\perp l; N\in BT, AN\perp BT;" O is the center of the circle ω.

"AF=7, AK=13, BT=11."

"\triangle AKF: KF=\sqrt{AK^2-AF^2}=\sqrt{13^2-7^2}=\sqrt{120}\approx 10.95."

"\angle AKF=arcsin\frac{AF}{AK}=arcsin\frac{7}{13}."

"\triangle AOK: \angle AOK=2\angle AKF=2arcsin\frac{7}{13}."

"OB=OA=OK=\frac{AK}{2sin\frac{\angle AOK}{2}}=\frac{13}{2\cdot \frac{7}{13}}=\frac{169}{14}."

"\triangle KBT: \angle BKT=arcsin\frac{BT}{BK}=arcsin\frac{11}{BK}."

"\triangle OBK: \angle BOK=2\angle BKT=2arcsin\frac{11}{BK}."

"BK=2OKsin\frac{\angle BOK}{2}=2\cdot \frac{ 169}{14}\cdot \frac{ 11}{BK},"

"BK=\sqrt{\frac{169\cdot11}{7}}\approx16.30."

"\triangle KBT: TK=\sqrt{BK^2-BT^2}=\sqrt{16.3^2-11^2}\approx 12.03."

"AN=TF=TK+KF=12.03+10.95=22.98."

"AF=NT=7, BN=BT-NT=11-7=4."

"\triangle ANB: AB=\sqrt{AN^2+BN^2}=\sqrt{22.98^2+4^2}\approx 23.33."

And by the Heron's formula, "S=102.35\approx 102."